# November 2025 On the Charm Contribution to the Muon g2g-2 Light-by-Light

### v1

@@ 1 Introduction
One of the most studied observables in particle physics in recent years is the anomalous magnetic moment of the muon Muong-2:2025xyk, Aoyama:2020ynm, Aliberti:2025beg. Much of the recent activity has focused on the hadronic light-by-light (HLbL) contribution, using phenomenological approaches Colangelo:2015ama, Masjuan:2017tvw, Colangelo:2017fiz, Hoferichter:2018kwz, Eichmann:2019tjk, Bijnens:2019ghy, Leutgeb:2019gbz, Cappiello:2019hwh, Masjuan:2020jsf, Bijnens:2020xnl, Bijnens:2021jqo, Danilkin:2021icn, Stamen:2022uqh, Leutgeb:2022lqw, Hoferichter:2023tgp, Hoferichter:2024fsj, Estrada:2024cfy, Ludtke:2024ase, Deineka:2024mzt, Eichmann:2024glq, Bijnens:2024jgh, Hoferichter:2024bae, Holz:2024diw, Cappiello:2025fyf and lattice QCD calculations Blum:2019ugy, Chao:2021tvp, Chao:2022xzg, Blum:2023vlm, Fodor:2024jyn. Although this piece is subleading, being suppressed by an additional factor $\mathcal{O}\!\left[(m_{\mu}/(2m_{c}))^{2}\right]$, there has also been interest in the charm-quark contribution in both communities, making it a useful benchmark for comparing QCD techniques. For the $\tau$ lepton, the analogous contribution is less suppressed and has recently been discussed in Ref. Hoferichter:2025fea.
@@ 2 Charm contribution to the light-by-light
The charm mass is known to be sufficiently large for perturbative QCD series to be well behaved in the first few orders. For Euclidean momenta, away from hadronic singularities, the charm mass acts as a regulator and allows for perturbative evaluations of Green functions even for $Q^{2}<m_{c}^{2}$. For example, see Figs. 7 and 8 of Ref. Davier:2023hhn for the two-point vector function. In this regime, the perturbative series take the form
@@
and power corrections are known to be suppressed in this regime Novikov:1977dq.
For the evaluation of the charm contribution to the light-by-light part of the muon g2g-2, the effective low-energy scale for the expansion is $m_{\mu}$. The lowest-order term was obtained long ago from the muon-loop contribution to the electron g2g-2. This calculation was performed fully analytically as an expansion in $m_{\mu}/m_{c}$ in Ref. Kuhn:2003pu, with the first two terms already given in Ref. Laporta:1992pa. These results were used in Refs. Aoyama:2020ynm, Aliberti:2025beg to estimate the charm contribution, assigning an uncertainty based on the bound $c\bar{c}$ states. The equal mass case is known fully analytically Laporta:1991zw.
@@
A representative lowest-order diagram is shown in Fig. 1(a). The evaluation to first order in $m_{\mu}/m_{c}$ is Laporta:1992pa, Boughezal:2011vw
@@
where $q_{c}=2/3$ is the charm-quark electric charge. The first correction in $\alpha_{s}$ comes from the type of diagrams shown in Fig. 1(b,c). These diagrams also enter the $\tau$-loop contribution to the muon g2g-2 and the muon-loop contribution to the electron g2g-2. The analytic expression to first order in $m_{\mu}/m_{c}$ was obtained in Ref. Boughezal:2011vw in a different context and is given by
@@
The result Section 2 of Ref. Boughezal:2011vw was obtained in the on-shell scheme for constituent-quark masses. We denote this mass by $M_{c}$. However, it is well known that the pole mass suffers from a renormalon ambiguity Beneke:1994sw, leading to poorly behaved perturbative series. This can be avoided by writing the expansions in terms of the $\overline{\mathrm{MS}}$ masses ParticleDataGroup:2024cfk. Indeed, it is known that typically expansions in the $\overline{\mathrm{MS}}$ scheme are known to converge better Maier:2007yn. To do this, we convert the pole mass to the $\overline{\mathrm{MS}}$ mass. The relation needed at this order was calculated long ago Tarrach:1980up, Gray:1990yh
@@
$A_{\mu}^{\text{HLbLc NLO, }\overline{\textrm{MS}}}$ $=A_{\mu}^{\text{HLbLc LO, }\overline{\textrm{MS}}}\dfrac{\alpha_{s}}{\pi}\left(C_{F}\dfrac{\Delta_{1}}{\Delta_{0}}-\dfrac{8}{3}+C\log\left(\dfrac{m_{c}(\mu)}{\mu}\right)\right)\,.$ (6)
@@
These values are taken from Ref. ParticleDataGroup:2024cfk and evolved down to $m_{c}$ using five-loop running, with a matching scale between $m_{b}$ and $2m_{b}$. The results are plotted in Fig. 2. We run $\alpha_{s}$ and $m_{c}$ from $m_{c}$ to $\mu$ at five-loop order with four active quark flavours, using both CRunDec Herren:2017osy and our own implementation, finding agreement. As can be seen, the LO+NLO result as a function of the subtraction scale $\mu$ is much better behaved.
Figure 2: The contribution from the charm quark to $a_{\mu}^{\textrm{HLbL}}$. LO $1/m_{c}^{10}$ is the result at lowest order including higher orders in $m_{\mu}/m_{c}$ Kuhn:2003pu. LO $1/m_{c}^{2}$ is the result of only the lowest order in $m_{\mu}/m_{c}$. LO+NLO is our full result with the central values of Section 2 as input, i.e. leading in $m_{\mu}/m_{c}$ but containing $\alpha_{s}$ corrections.
@@
This corresponds to the black line with shaded background in Fig. 2 and covers the entire area shown in Fig. 3. This represents a significantly improved error compared to the result used in Aoyama:2020ynm, Aliberti:2025beg of $3(1)\times 10^{-11}$ in the analytical and dispersive section.
Knowledge of higher orders in these perturbative series would help to further reduce the associated uncertainties. While these terms are not known, we can support our uncertainty estimates by applying the same criterion to other RGE-invariant series for which higher-order coefficients are available Maier:2007yn, Maier:2009fz. The first moment of the non-singlet vector two-point function is known up to four loops and, including subsequent perturbative orders, one finds,
@@
$=3.29\times 10^{-11}~(\mathrm{LO}+\mathrm{NLO})\,.$ (10)
The labels have the same meaning as in Fig. 2. Here one notices that we obtain a very large $\alpha_{s}$ correction, much larger than in Fig. 2. Assuming that higher-order corrections are also large, this is in reasonable agreement with Eq. 8.
@@
The tau can be treated in the same way. This has been evaluated in Ref. Hoferichter:2025fea for the full quark-loop at $m_{c}(m_{c})$ with $a_{\tau}^{\textrm{HLbLc}}=4.5\times 10^{-9}$. Our results are at the leading order in $1/m_{c}^{2}$ and are shown in Fig. 5.
@@ 3 Final discussion
We have incorporated the $\alpha_{s}$ correction to the charm-loop contribution to the muon HLbL g2g-2 in the $\overline{\mathrm{MS}}$ scheme, translating the $\alpha_{s}$ corrections to (model-dependent) constituent quark contributions computed in Ref. Boughezal:2011vw. We have argued that this leads to better-controlled perturbative uncertainties. Our final number is
@@
which can be compared to $a_{\mu}^{\mathrm{HLbLc}}=3(1)\times 10^{-11}$ from Ref. Aliberti:2025beg, obtained without incorporating gluon corrections, or to the slightly different value of Ref. Hoferichter:2025fea, $a_{\mu}^{\mathrm{HLbLc}}=2.7(8)\times 10^{-11}$, which attempts to consider corrections to the quark loop with a different method.
The charm quark contribution has also been calculated using lattice QCD. There are two full calculations. The Mainz/CLS collaboration finds Chao:2022xzg
@@
The final result of the BMW collaboration is Fodor:2024jyn
@@
These calculations also include estimates of the disconnected contribution. This does not appear in our result, since it starts contributing only at order $\alpha_{s}^{2}$, beyond the order we have considered. This contribution was calculated by BMW and found to be 0.185(54)-0.185(54). The BMWc result is in very good agreement with ours, while there is mild tension with the Mainz/CLS result.

### v2

# December 2025 On the Charm Contribution to the Muon g2g-2 Light-by-Light

@@ 1 Introduction
One of the most studied observables in particle physics in recent years is the anomalous magnetic moment of the muon Aguillard and others [2025], Aoyama and others [2020], Aliberti and others [2025]. Much of the recent activity has focused on the hadronic light-by-light (HLbL) contribution, using phenomenological approaches Colangelo et al. [2015], Masjuan and Sánchez-Puertas [2017], Colangelo et al. [2017], Hoferichter et al. [2018], Eichmann et al. [2019], Bijnens et al. [2019], Leutgeb and Rebhan [2020], Cappiello et al. [2020], Masjuan et al. [2022], Bijnens et al. [2020, 2021], Danilkin et al. [2021], Stamen et al. [2022], Leutgeb et al. [2023], Hoferichter et al. [2023, 2024], Estrada et al. [2024], Lüdtke et al. [2025], Deineka et al. [2025], Eichmann et al. [2025], Bijnens et al. [2025], Hoferichter et al. [2025a], Holz et al. [2025], Cappiello et al. [2025] and lattice QCD calculations Blum et al. [2020], Chao et al. [2021, 2022], Blum et al. [2025], Fodor et al. [2025]. Although this piece is subleading, being suppressed by an additional factor $\mathcal{O}\!\left[(m_{\mu}/(2m_{c}))^{2}\right]$, there has also been interest in the charm-quark contribution in both communities, making it a useful benchmark for comparing QCD techniques. For the $\tau$ lepton, the analogous contribution is less suppressed and has recently been discussed in Ref. Hoferichter et al. [2025b].
@@ 2 Charm contribution to the light-by-light
The charm mass is known to be sufficiently large for perturbative QCD series to be well behaved in the first few orders. For Euclidean momenta, away from hadronic singularities, the charm mass acts as a regulator and allows for perturbative evaluations of Green functions even for $Q^{2}<m_{c}^{2}$. For example, see Figs. 7 and 8 of Ref. Davier et al. [2023] for the two-point vector function. In this regime, the perturbative series take the form
@@
and power corrections are known to be suppressed in this regime Novikov et al. [1978].
For the evaluation of the charm contribution to the light-by-light part of the muon g2g-2, the effective low-energy scale for the expansion is $m_{\mu}$. The lowest-order term was obtained long ago from the muon-loop contribution to the electron g2g-2. This calculation was performed fully analytically as an expansion in $m_{\mu}/m_{c}$ in Ref. Kuhn et al. [2003], with the first two terms already given in Ref. Laporta and Remiddi [1993]. These results were used in Refs. Aoyama and others [2020], Aliberti and others [2025], Colangelo et al. [2020] to estimate the charm contribution, assigning an uncertainty based on the bound $c\bar{c}$ states. The equal mass case is known fully analytically Laporta and Remiddi [1991].
@@
A representative lowest-order diagram is shown in Fig. 1(a). The evaluation to first order in $m_{\mu}/m_{c}$ is Laporta and Remiddi [1993], Boughezal and Melnikov [2011]
@@
where $q_{c}=2/3$ is the charm-quark electric charge. The first correction in $\alpha_{s}$ comes from the type of diagrams shown in Fig. 1(b,c). These diagrams also enter the $\tau$-loop contribution to the muon g2g-2 and the muon-loop contribution to the electron g2g-2. The analytic expression to first order in $m_{\mu}/m_{c}$ was obtained in Ref. Boughezal and Melnikov [2011] in a different context and is given by
@@
The result Section 2 of Ref. Boughezal and Melnikov [2011] was obtained in the on-shell scheme for constituent-quark masses. We denote this mass by $M_{c}$. However, it is well known that the pole mass suffers from a renormalon ambiguity Beneke and Braun [1994], leading to poorly behaved perturbative series. This can be avoided by writing the expansions in terms of the $\overline{\mathrm{MS}}$ masses Navas and others [2024]. Indeed, it is known that typically expansions in the $\overline{\mathrm{MS}}$ scheme are known to converge better Maier et al. [2008]. To do this, we convert the pole mass to the $\overline{\mathrm{MS}}$ mass. The relation needed at this order was calculated long ago Tarrach [1981], Gray et al. [1990]
@@
$A_{\mu}^{\text{HLbLc NLO, }\overline{\textrm{MS}}}$ $=A_{\mu}^{\text{HLbLc LO, }\overline{\textrm{MS}}}\dfrac{\alpha_{s}}{\pi}$
$~~~~~\times\left(C_{F}\dfrac{\Delta_{1}}{\Delta_{0}}-\dfrac{8}{3}+C\log\left(\dfrac{m_{c}(\mu)}{\mu}\right)\right)\,.$ (6)
@@
These values are taken from Ref. Navas and others [2024] and evolved down to $m_{c}$ using five-loop running, with a matching scale between $m_{b}$ and $2m_{b}$. The results are plotted in Fig. 2. We run $\alpha_{s}$ and $m_{c}$ from $m_{c}$ to $\mu$ at five-loop order with four active quark flavours, using both CRunDec Herren and Steinhauser [2018] and our own implementation, finding agreement. As can be seen, the LO $1/m_{c}^{2}$+NLO $1/m_{c}^{2}$ result as a function of the subtraction scale $\mu$ is much better behaved. Using instead LO $1/m_{c}^{10}$+NLO $1/m_{c}^{2}$ does not change the conclusion.
[…]
@@
This corresponds to the black line with shaded background in Fig. 2 and covers the entire area shown in Fig. 3. This represents a significantly improved error compared to the result used in Aoyama and others [2020], Aliberti and others [2025], Colangelo et al. [2020] of $3(1)\times 10^{-11}$ in the analytical and dispersive section.
Knowledge of higher orders in these perturbative series would help to further reduce the associated uncertainties. While these terms are not known, we can support our uncertainty estimates by applying the same criterion to other RGE-invariant series for which higher-order coefficients are available Maier et al. [2008, 2010]. The first moment of the non-singlet vector two-point function is known up to four loops and, including subsequent perturbative orders, one finds,
@@
$=3.29\times 10^{-11}~(\mathrm{LO}~1/M_{c}^{2}+\mathrm{NLO}~1/M_{c}^{2})\,.$ (10)
The labels have the same meaning as in Fig. 2. Here one notices that we obtain a very large $\alpha_{s}$ correction, much larger than in Fig. 2. Assuming that higher-order corrections are also large, [^1] For comparison, the equivalent of Eq. 9 in the on-shell scheme, taking $\alpha_{s}(M_{c})=0.3519$, is $\Pi_{V}(Q\ll M_{c})\propto(1+0.567+0.391+0.483)\,z(M_{c})$, thus having very large higher order corrections. this is in reasonable agreement with Eq. 8.
@@
The tau can be treated in the same way. This has been evaluated in Ref. Hoferichter et al. [2025b] for the full quark-loop at $m_{c}(m_{c})$ with $a_{\tau}^{\textrm{HLbLc}}=4.5\times 10^{-9}$. Our results are at the leading order in $1/m_{c}^{2}$ and are shown in Fig. 5.
@@ 3 Final discussion
We have incorporated the $\alpha_{s}$ correction to the charm-loop contribution to the muon HLbL g2g-2 in the $\overline{\mathrm{MS}}$ scheme, translating the $\alpha_{s}$ corrections to (model-dependent) constituent quark contributions computed in Ref. Boughezal and Melnikov [2011]. We have argued that this leads to better-controlled perturbative uncertainties. Our final number is
@@
which can be compared to $a_{\mu}^{\mathrm{HLbLc}}=3(1)\times 10^{-11}$ from Ref. Aliberti and others [2025], obtained without incorporating gluon corrections, or to the slightly different value of Ref. Hoferichter et al. [2025b], $a_{\mu}^{\mathrm{HLbLc}}=2.7(8)\times 10^{-11}$, which attempts to consider corrections to the quark loop with a different method.
The charm quark contribution has also been calculated using lattice QCD. There are two full calculations. The Mainz/CLS collaboration finds Chao et al. [2022]
@@
The final result of the BMW collaboration is Fodor et al. [2025]
@@
The BMWc result is in very good agreement with ours, while there is mild tension with the Mainz/CLS result. These calculations also include estimates of the disconnected contribution. This does not appear in our result, since it starts contributing only at order $\alpha_{s}^{2}$, beyond the order we have considered, and also contains a nonperturbative part. This contribution was calculated by BMW and found to be 0.185(52)(04)-0.185(52)(04) consistent with the Mainz/CLS calculation. As this is a sizable shift, we can compare to the connected lattice QCD charm-quark result in isolation. The connected part of the BMWc the result is Fodor et al. [2025]
$a_{\mu}^{\textrm{HLbLc, conn.}}=3.92(26)\times 10^{-11}\,.$ (15)
The Mainz/CLS collaboration has Chao et al. [2022]
$a_{\mu}^{\textrm{HLbLc, conn.}}=3.1(4)\times 10^{-11}\,.$ (16)
As can be seen, our value is consistent within errors.